# Scalaron dark matter dynamics: effects of Higgs non-minimal coupling to gravity

### v1

@@ Abstract
One of the key features of the $R^{2}$-gravity is the embedding of a scalar field, scalaron, into the gravity sector. The scalaron interacts with the Standard Model (SM) matter fields through Planck-suppressed couplings. If the scalaron serves as a viable dark matter (DM) candidate, it can account for the lack of evidence of DM interactions beyond gravity in experimental and observational probes to date. The realization of the scalaron, as a cold DM candidate, depends on an induced trilinear interaction with the SM Higgs, despite its suppressed coupling strength. Here, we introduce a Higgs non-minimal coupling to gravity that additionally contributes to the induced trilinear interaction with its existing competing part, originated from the $R^{2}$-gravity. We study the interplay between these two contributions in the early universe, which determines both the initial conditions and evolution of the scalaron, leading to cold DM behavior at a later epoch. The trilinear interaction vanish identically for certain combinations of the Higgs non-minimal coupling and the scalaron mass, thereby setting the scalaron density through misalignment mechanism, similar to axions. Consequently, we find that the observed DM relic density is obtained for a scalaron mass between $1.38$ keV and $0.7$ MeV. The lower limit on the mass stems from the LHC constraints on the Higgs non-minimal coupling, whereas the upper bound arises from INTEGRAL/SPI limits on the excess gamma-ray flux from possible scalaron decays to two photons.
@@ 1 Introduction
As a nightmare scenario, DM may not interact with SM particles at all, which does not hinder the gravitational clustering, essential for the LSS formation. It only depends on the DM density during the matter-radiation equality epoch, in the cold DM paradigm. Then, the question comes why there is little or absolutely no interaction between DM and SM particles. In this context, extensions of Einstein’s general relativity (GR), particularly $R^{2}$-gravity ($R$ is the Ricci scalar) can provide an explanation to naturally suppressed interactions of DM with the rest of matter fields Cembranos:2008gj ; Corda:2011aa ; Katsuragawa:2016yir ; Katsuragawa:2017wge ; Shtanov:2021uif . In this case, an extra scalar degree of freedom is hidden inside the gravity sector, in addition to two polarizations of gravitational waves. The scalar field becomes disentangled from the gravity sector, while restoring the usual Einstein-Hilbert action DeFelice:2010aj , through a conformal transformation of the original theory with $R^{2}$ term. Since the scalar field, often termed as scalaron starobinsky , comes from the gravity sector, it couples democratically to all other preexisting matter fields but with Planck-scale suppressed couplings after the transformation. Therefore, the scalaron naturally explains the non-observation of DM-SM interactions in existing experiments till date, subject to its viability as DM candidate.
@@
How can the scalaron density satisfy the observed relic density of DM ? The scalar field dynamics, leading to the present-day DM density is a well studied scenario in the context of axion cosmology Preskill:1982cy ; Abbott:1982af ; Dine:1982ah ; Marsh:2015xka . The axion field is misaligned initially from its potential minima. When the Hubble friction becomes negligible compared to the axion mass, coherent oscillation (i.e. associated quanta are in the zero momentum state) starts about the minima of its potential, approximated to a quadratic form. The energy density of the axion field during coherent oscillation redshifts as pressure-less matter, thereby validating it as a cold DM candidate. The observed DM relic density determines the axion mass, $m\sim\mathcal{O}(\mu\rm eV)$ and the initial field value, $\phi_{i}\sim 10^{10}{~\rm GeV}$, for the misalignment scenario during a radiation-dominated universe.
@@
The misalignment mechanism in the context of scalaron has been discussed in Ref.Cembranos:2008gj , without taking effects of scalaron interactions with the SM particles in determining its dynamics in the early universe. An induced trilinear interaction of the scalaron with the Higgs, fixed by the Higgs quartic self-coupling and the vacuum expectation value (VEV), determines the initial condition and the subsequent scalaron dynamics, despite the Planck suppression stemming from its gravitational origin. Thus, the evolution of the scalaron is connected to the Electroweak phase transition (EWPT) Shtanov:2021uif ; Shtanov:2022xew , during which the Higgs field moves from a trivial minima ($\langle h\rangle=0$) to its SM VEV. In this scenario, the coherent oscillation of the scalaron starts only when the Higgs field acquires its SM VEV, unlike the former case, where the Hubble friction decides the onset of the oscillation. As a result, the scalaron saturates the DM density only for a very restrictive mass around $\mathcal{O}$(meV). In contrast, the scalaron mass for the case in Ref.Cembranos:2008gj , varies from $\mathcal{O}$(meV) to $\mathcal{O}$(MeV), depending on the initial field value, which is set independently of Higgs dynamics during the EWPT.
@@
In the present work, we introduce Higgs non-minimal coupling to gravity in addition to the $R^{2}$ term in the Einstein-Hilbert action, for which the above-mentioned scenarios can be realized in a general framework. In particular, the trilinear scalaron-Higgs interaction now gets competing contributions from both $R^{2}$ term and the Higgs non-minimal coupling to gravity. We find that an interplay between these terms decides the initial condition and the scalaron dynamics in the early universe. In fact, the interaction can identically vanish for different combinations of non-minimal coupling and the scalaron mass. Consequently, the scalaron density is set via misalignment dynamics, as in the axion cosmology. In this case, we find that the scalaron mass varies from $\mathcal{O}$(keV) nearly to $\mathcal{O}$(MeV), while satisfying the DM relic density constraint, Higgs global signal strength measurement at Large Hadron Collider (LHC) and $16$ years of gamma-ray observations by INTEGRAL/SPI telescope. For a non-zero scalaron-Higgs interaction, we find that the scalaron mass is constrained to $\mathcal{O}$(meV), irrespective of Higgs non-minimal coupling to gravity permitted by LHC constraints.
@@
The organization of the paper is as follows : the general framework for scalaron dynamics in presence of the Higgs non-minimal coupling has been discussed in Sec.2, followed by the detailed analytic calculation of scalaron dynamics in Sec.3. Then we discuss observational constraints on the non-minimal coupling from LHC experiments and the gamma ray constraints on scalaron decays to two photons from the INTEGRAL/SPI telescope in Sec.4. Finally we summarize our findings in Sec.5.
@@ 2 The general framework
We first write the Einstein-Hilbert action with an additional $R^{2}$ term in the Jordan frame as the following starobinsky ; Shtanov:2021uif :
@@
where $M_{P}(=\sqrt{3/16\pi G}\approx 3\times 10^{18}$ GeV) is the normalized Planck mass. [^1] We have followed the notation of Shtanov:2021uif for the normalized Planck mass $M_{P}$ so that the conformal factor comes out to be $e^{-\frac{\phi}{M_{p}}}$, instead of $e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}$. We do not include the cosmological constant in the gravity lagrangian, since our present study is related to a standard radiation-dominated universe only. At this point, we note that the scale $m$ is going to determine the mass of the scalar field responsible for the DM, and it is in the range of $\rm meV$ to $\rm MeV$.
@@
where $\Phi$ is the Higgs doublet, $v~(=246~{\rm GeV})$ is the Higgs VEV at zero temperature and $\lambda$ is the Higgs quartic self-coupling with its SM value, $\lambda\approx 0.13$. We shall demonstrate that the current theory with Eq. (1) and Eq. (2) will generate an interaction between the scalaron and the Higgs at the renormalizable level. In the context of inflation, such theories with both $R^{2}$-gravity and Higgs non-minimal coupling terms have been considered in the literature Ema:2017rqn ; Gorbunov:2018llf ; He:2018mgb ; He:2018gyf ; Gundhi:2018wyz ; Cheong:2021vdb .
To note, the non-minimal coupling, $\xi$ is a dimensionless quantity, which couples Higgs to gravity at the renormalizable level. This coupling can be constrained (detailed in Sec.4) from the measurement of Higgs global signal strength at LHC, given by $|\xi|<1.38\times 10^{15}$ at $95\%$ confidence level (C.L.). The sign of $\xi$ can not be determined uniquely from the signal strength, since only the magnitude of $\xi$ appears in the amplitudes of physical processes, measured at the LHC. However, to ensure the vacuum stability of Higgs potential during high scale inflationary epoch, $\xi$ turns out to be a positive quantity, particularly $\xi\gtrsim 0.06$ Herranen:2014cua . Also for post-inflationary epoch $\xi$ is reported to be positive Figueroa:2017slm . Although we are not concerned with inflationary dynamics in the present context, we work with positive values of $\xi$, i.e. $0\leq\xi\leq 1.38\times 10^{15}$.
To express the total action, (i.e. $S=S_{g}+S_{M}$) in the Einstein frame, we first introduce an auxiliary field, $\sigma$ and write an equivalent action (well known technique in $f(R)$ gravity theories, see for example Refs.Faraoni:2004pi ; DeFelice:2010aj ) as follows :
$S=-\frac{M^{2}_{P}}{3}\int d^{4}x\sqrt{-g}\left[\left(1-\frac{\sigma}{3m^{2}}+\frac{3\xi}{M^{2}_{P}}\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)\right)R+\frac{\sigma}{6m^{2}}\right]$ (3)
@@
$V(\phi,\tilde{\Phi})\simeq\frac{1}{2}m^{2}M_{P}^{2}\left[1-e^{-\phi/M_{P}}-\frac{3\xi}{M_{P}^{2}}\left(\tilde{\Phi}^{\dagger}\tilde{\Phi}-e^{-\phi/M_{P}}\frac{v^{2}}{2}\right)\right]^{2}+\lambda\left(\tilde{\Phi}^{\dagger}\tilde{\Phi}-e^{-\phi/M_{P}}\frac{v^{2}}{2}\right)^{2},$ (6)
while dropping non-renormalizable terms suppressed by the Planck mass. For sub-Planckian value of $\phi$, i.e. $\phi/M_{P}\ll 1$ and taking the Higgs doublet in the unitary gauge as $\tilde{\Phi}^{T}=(0~~h/\sqrt{2})$, the potential with dominant contributions becomes,
@@
It is apparent from the above conditions that for non-zero trilinear interaction with the Higgs, the minima of the scalaron becomes non-trivial before the EWPT ; subsequently the minima shifts towards zero as $T\ll T_{c}$. Hence, the initial minima of the scalaron determines the energy density of the coherent oscillation of the field around its final minima, i.e. $\phi_{0}=0$. This is a quite departure from the standard misalignment scenario, where the initial condition is usually set away from the minima, which provides the energy of the coherent oscillation. In the present situation, the source of that energy comes from the interaction with the Higgs, especially due to Higgs quartic self-coupling. Thus the scalaron, being in its natural minima at the early epoch, can account for the DM density at a later epoch.
The interaction with Higgs vanishes when $3\xi m^{2}=\lambda v^{2}$, in presence of a non-zero non-minimal coupling of Higgs to gravity. Consequently, the scalaron minima restores to $\phi=0$ for all temperatures, signifying complete decoupling from the Higgs sector from the very beginning. In this case, the DM density is satisfied by the scalaron, if it is misaligned from its minima in the early universe, reminiscent of the axion cosmology.
@@
As we solve the equation in the radiation-dominated universe from the epoch of the EWPT ($T=T_{c}$), the time derivatives are conveniently transformed into temperature derivatives, using the relation, $\dot{T}=-HT$, where $H=(\frac{\pi^{2}}{60}g_{*})^{1/2}\frac{T^{2}}{M_{P}}$. Here, we neglect the variation of relativistic degrees of freedom ($g_{*}$). We write the EOM in terms of rescaled dimensionless variables, $\tilde{\phi}=\frac{\phi}{\phi_{0}(T_{c})}$ (for $\phi_{0}\neq 0$) and $\tilde{T}=\frac{T}{T_{c}}$, i.e.
@@
but with an arbitrary initial condition away from its minima, i.e. $\phi=\phi_{i}$ at $T=T_{i}$, for which we define $\bar{\phi}=\phi/\phi_{i}$ and $\bar{T}=T/T_{i}$. Notably, the initial temperature in this scenario can be arbitrary; whereas in the former one, the initial temperature is the critical temperature of the EWPT. This is due to apparent decoupling of the scalaron from the Higgs sector, leaving scalaron dynamics unaffected by the phase transition.

### v2

@@ Abstract
One of the key features of the $R^{2}$-gravity is the embedding of a scalar field, scalaron, into the gravity sector. The scalaron interacts with the Standard Model (SM) matter fields through Planck-suppressed couplings. If the scalaron serves as a viable dark matter (DM) candidate, it can account for the lack of evidence of DM interactions beyond gravity in experimental and observational probes to date. The realization of the scalaron, as a cold DM candidate, depends on an induced trilinear interaction with the SM Higgs via its quartic self coupling. Here, we introduce a Higgs non-minimal coupling to gravity that additionally contributes to the induced trilinear interaction with its existing competing part, originated from the $R^{2}$-gravity. We study the interplay between these two contributions in the early universe, which determines both the initial conditions and evolution of the scalaron, leading to cold DM behavior at a later epoch. The trilinear interaction vanishes at the leading order for certain combinations of the Higgs non-minimal coupling ($\xi$) and the scalaron mass ($m$), thereby setting the scalaron density through misalignment mechanism, similar to axions. In this case, the scalaron DM mass is obtained as, $2.7~{\rm meV}\lesssim m\lesssim 0.7~\rm{MeV}$. The lower limit on the mass is set by the fifth force constraints from torsion balance experiment, whereas the upper bound arises from INTEGRAL/SPI limits on the excess gamma-ray flux due to possible decay of scalaron into two photons. On the other hand, when the trilinear interaction is non-zero and dominated by the Higgs quartic self coupling, the DM relic density is satisfied with $m\simeq 3.6$ meV. Alternately, when the Higgs non-minimal coupling to gravity dominates, the mass is found to be in the range 10–120 meV. We also obtain, for the first time within the scalaron-Higgs mixed model, an upper bound on $|\xi m|$ of $2.35\times 10^{16}$ GeV, from Higgs mass measurement at the LHC.
@@ 1 Introduction
As a nightmare scenario, DM may not interact with SM particles at all, which does not hinder the gravitational clustering, essential for the LSS formation. It only depends on the DM density during the matter-radiation equality epoch, in the cold DM paradigm. Then, the question comes why there is little or absolutely no interaction between DM and SM particles. In this context, extensions of Einstein’s general relativity (GR), particularly $R^{2}$-gravity ($R$ is the Ricci scalar) can provide an explanation to naturally suppressed interactions of DM with the rest of matter fields Cembranos:2008gj ; Corda:2011aa ; Katsuragawa:2016yir ; Katsuragawa:2017wge ; Shtanov:2021uif . In this case, an extra scalar degree of freedom is hidden inside the gravity sector, in addition to two polarizations of gravitational waves. The scalar field becomes disentangled from the gravity sector, while restoring the usual Einstein-Hilbert action Nojiri:2017ncd ; DeFelice:2010aj , through a conformal transformation of the original theory with $R^{2}$ term. Since the scalar field, often termed as scalaron starobinsky ; Nojiri:2008nt , comes from the gravity sector, it couples democratically to all other preexisting matter fields but with Planck-scale suppressed couplings after the transformation. Therefore, the scalaron naturally explains the non-observation of DM-SM interactions in existing experiments till date, subject to its viability as DM candidate.
@@
How can the scalaron density satisfy the observed relic density of DM ? The scalar field dynamics, leading to the present-day DM density is a well studied scenario in the context of axion cosmology Preskill:1982cy ; Abbott:1982af ; Dine:1982ah ; Marsh:2015xka . The axion field is misaligned initially from its potential minima. When the Hubble friction becomes negligible compared to the axion mass, coherent oscillation (i.e., associated quanta are in the zero momentum state) starts about the minima of its potential, approximated to a quadratic form. The energy density of the axion field during coherent oscillation redshifts as pressure-less matter, thereby validating it as a cold DM candidate. The observed DM relic density determines the axion mass, $m\sim\mathcal{O}(\mu\rm eV)$ and the initial field value, $\phi_{i}\sim 10^{10}{~\rm GeV}$, for the misalignment scenario during a radiation-dominated universe.
@@
The misalignment mechanism in the context of scalaron has been discussed in Ref.Cembranos:2008gj , without taking effects of scalaron interactions with the SM particles in determining its dynamics in the early universe. An induced trilinear interaction of the scalaron with the Higgs, fixed by the Higgs quartic self-coupling and the vacuum expectation value (VEV), determines the initial condition and the subsequent scalaron dynamics, despite the Planck suppression stemming from its gravitational origin. Thus, the evolution of the scalaron is connected to the Electroweak phase transition (EWPT) Shtanov:2021uif ; Shtanov:2022xew , during which the Higgs field moves from a trivial minima ($\langle h\rangle=0$) to its SM VEV. In this scenario, the coherent oscillation of the scalaron starts only when the Higgs field acquires its SM VEV, unlike the former case, where the Hubble friction decides the onset of the oscillation. As a result, the scalaron saturates the DM density only for a very restrictive mass around $\mathcal{O}$(meV). In contrast, the scalaron mass for the case in Ref.Cembranos:2008gj , varies from $\mathcal{O}(\text{meV})$ to $\mathcal{O}(\text{MeV})$, depending on the initial field value, which is set independently of Higgs dynamics during the EWPT.
@@
In the present work, we introduce Higgs non-minimal coupling to gravity ($\xi$) in addition to the $R^{2}$ term in the Einstein-Hilbert action, for which the above-mentioned scenarios can be realized in a general framework. In particular, the trilinear scalaron-Higgs interaction now gets competing contributions from both $R^{2}$ term and the Higgs non-minimal coupling to gravity. We find that an interplay between these terms decides the initial condition and the scalaron dynamics in the early universe. When the interaction is dominated by the Higgs quartic coupling, the scalaron mass around a few meV fixes the DM relic density. On the other hand, the scalaron mass can vary from $\mathcal{O}(10-100)$ meV in case of Higgs non-minimal coupling to gravity dominating the interaction. The upper limit on the scalaron mass in this scenario comes from the Higgs mass measurement at the Large hadron Collider (LHC).
The trilinear interaction vanishes at the leading order for certain combinations of $\xi$ and the scalaron mass, $m$. Consequently, the scalaron density is set via misalignment dynamics, as in the axion cosmology. In this case, we find that the DM relic density is satisfied for a range of scalaron masses, i.e., from $\mathcal{O}$(meV) nearly to $\mathcal{O}$(MeV). This lower limit on $m$ comes from the fifth force constraint obtained in the precision torsion balance experiment, while the upper bound stems from $16$ years of gamma-ray observations by INTEGRAL/SPI telescope.
@@
The organization of the paper is as follows : the general framework for scalaron dynamics in presence of the Higgs non-minimal coupling is discussed in Sec.2, followed by the detailed analytic calculation of scalaron dynamics in Sec.3. In Sec. 4, we discuss the relevant experimental constraints on the scenarios considered, in particular, the LHC constraints on $|\xi m|$ in our scalaron-Higgs mixed model, followed by the gamma ray constraints on possible decay of the scalaron into a pair of photons from the INTEGRAL/SPI telescope and the fifth force constraint on the scalaron mass. Finally, we summarize our findings in Sec. 5 and present our concluding remarks in Sec. 6.
@@ 2 The general framework
We first write the Einstein-Hilbert action with an additional $R^{2}$ term in the Jordan frame as the following starobinsky ; Shtanov:2021uif ; Oikonomou:2025htz :
@@
where $M_{P}(=\sqrt{3/16\pi G}\approx 3\times 10^{18}$ GeV) is the normalized Planck mass. [^1] We have followed the notation of Shtanov:2021uif for the normalized Planck mass $M_{P}$ so that the conformal factor comes out to be $e^{-\frac{\phi}{M_{P}}}$, instead of $e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}$. We do not include the cosmological constant in the gravity lagrangian, since our present study is related to a standard radiation-dominated universe only. At this point, we note that the scale $m$ is going to determine the mass of the scalar field responsible for the DM, and it is in the range of meV to $\rm MeV$.
@@
where $\Phi$ is the Higgs doublet, $v~(=246~{\rm GeV})$ is the Higgs VEV at zero temperature and $\lambda$ is the Higgs quartic self-coupling with its SM value, $\lambda\approx 0.13$. We shall demonstrate that the current theory with Eq. (1) and Eq. (2) will generate an interaction between the scalaron and the Higgs at the renormalizable level. In the context of inflation, such theories with both $R^{2}$-gravity and Higgs non-minimal coupling terms have been considered in the literature Ema:2017rqn ; Gorbunov:2018llf ; He:2018mgb ; He:2018gyf ; Gundhi:2018wyz ; Cheong:2021vdb ; Canko:2019mud ; Wang:2017fuy ; Salvio:2015kka ; Salvio:2016vxi .
To note, the non-minimal coupling, $\xi$ is a dimensionless quantity, which couples Higgs to gravity at the renormalizable level, thereby potentially can modify the Higgs mass and its decay widths. This coupling solely can not be constrained from the Higgs signal strength at LHC within the scalaron-Higgs mixed model, rather we obtain an upper bound on $|\xi m|$ to be $2.35\times 10^{16}$ GeV at $95\%$ confidence level (C.L.). For the detailed derivation of this result see Sec.4.1. The sign of $\xi$ can not be determined uniquely from the signal strength, since only the magnitude of $\xi$ appears in the amplitudes of physical processes, measured at the LHC. However, to ensure the vacuum stability of Higgs potential during high scale inflationary epoch, $\xi$ turns out to be a positive quantity, particularly $\xi\gtrsim 0.06$ Herranen:2014cua . Also for post-inflationary epoch $\xi$ is reported to be positive Figueroa:2017slm . Although we are not concerned with inflationary dynamics in the present context, we work with positive values of $\xi$, i.e., $0\leq\xi m\leq 2.35\times 10^{16}$ GeV.
To express the total action, (i.e., $S=S_{g}+S_{M}$) in the Einstein frame, we first introduce an auxiliary field, $\sigma$ and write an equivalent action (well known technique in $f(R)$ gravity theories, see for example Refs.Faraoni:2004pi ; DeFelice:2010aj ) as follows :
$S=-\frac{M^{2}_{P}}{3}\int d^{4}x\sqrt{-g}\left[\left(1-\frac{\sigma}{3m^{2}}+\frac{3\xi}{M^{2}_{P}}\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)\right)R+\frac{\sigma^{2}}{6m^{2}}\right]$ (3)
@@
$V(\phi,\tilde{\Phi})=\frac{1}{2}m^{2}M_{P}^{2}\left[1-e^{-\phi/M_{P}}-\frac{3\xi}{M_{P}^{2}}\left(\tilde{\Phi}^{\dagger}\tilde{\Phi}-e^{-\phi/M_{P}}\frac{v^{2}}{2}\right)\right]^{2}+\lambda\left(\tilde{\Phi}^{\dagger}\tilde{\Phi}-e^{-\phi/M_{P}}\frac{v^{2}}{2}\right)^{2},$ (6)
while dropping non-renormalizable terms suppressed by the Planck mass. For sub-Planckian value of $\phi$, i.e., $\phi/M_{P}\ll 1$ and taking the Higgs doublet in the unitary gauge as $\tilde{\Phi}^{T}=(0~~h/\sqrt{2})$, the potential with dominant contributions becomes,
@@
It is apparent from the above conditions that for non-zero trilinear interaction with the Higgs, the minima of the scalaron becomes non-trivial before the EWPT ; subsequently the minima shifts towards zero as $T\ll T_{c}$. Hence, the initial minima of the scalaron determines the energy density of the coherent oscillation of the field around its final minima, i.e., $\phi_{0}=0$. This is a quite departure from the standard misalignment scenario, where the initial condition is usually set away from the minima, which provides the energy of the coherent oscillation. In the present situation, the source of that energy comes from the interaction with the Higgs via $\lambda$ and $\xi$. Thus the scalaron, being in its natural minima at the early epoch, can account for the DM density at a later epoch.
In presence of Higgs non-minimal coupling to gravity the trilinear interaction vanish at the leading order for $3\xi m^{2}=\lambda v^{2}$. Consequently, we can approximate the scalaron evolution to be decoupled from the Higgs sector from the very beginning and the scalaron minima to be $\phi\approx 0$ for all temperatures. In passing, we note that the interaction does not identically vanish under the present condition, rather leaves a residual term, suppressed by $1/M^{3}_{P}$. However, it leads to a small initial field value, $\frac{\phi_{0}(T_{c})}{M_{P}}=\frac{\lambda^{2}v^{8}}{4m^{4}M_{p}^{4}}\sim 10^{-12}$, thereby effectively needing an extra initial condition for the scalaron to be a valid DM candidate, as shown in the subsequent study.
@@
As we solve the equation in the radiation-dominated universe from the epoch of the EWPT ($T=T_{c}$), the time derivatives are conveniently transformed into temperature derivatives, using the relation, $\dot{T}=-HT$, where $H=(\frac{\pi^{2}}{60}g_{*})^{1/2}\frac{T^{2}}{M_{P}}$. Here, we neglect the variation of relativistic degrees of freedom ($g_{*}$). We write the EOM in terms of rescaled dimensionless variables, $\tilde{\phi}=\frac{\phi}{\phi_{0}(T_{c})}$ (for $\phi_{0}\neq 0$) and $\tilde{T}=\frac{T}{T_{c}}$, i.e.,
@@
but with an arbitrary initial condition away from its minima, i.e., $\phi=\phi_{i}$ at $T=T_{i}$, for which we define $\bar{\phi}=\phi/\phi_{i}$ and $\bar{T}=T/T_{i}$. Notably, the initial temperature in this scenario can be arbitrary; whereas in the former one, the initial temperature is the critical temperature of the EWPT. This is due to apparent decoupling of the scalaron from the Higgs sector, leaving scalaron dynamics unaffected by the phase transition.